a wedged-shaped piece is cut from a circular pizza. the radius of the pizza is 6 inches. the rounded edge of the cust of the piece measuures 4.2 inches. to the nearest tenth the angle of the pointed end of the pizza in radians is ?

Answer by lwsshak3(11628)

a wedged-shaped piece is cut from a circular pizza. the radius of the pizza is 6 inches. the rounded edge of the cust of the piece measuures 4.2 inches. to the nearest tenth the angle of the pointed end of the pizza in radians is ?
**
circumference of pizza=2πr
entire circumference covers 2π radians
(4.2/2πr)*2π=4.2/r=4.2/6=0.7 radians